OpenStudy (anonymous):

Melissa wants to spend no more than $225 on school clothes. She spends $100 on a coat and then wants to buy some sweaters that are on special for $30 each. Solve the inequality 100 + 30s ≤ 225 to find the greatest number of sweaters s she can buy.

OpenStudy (anonymous):

100+30s,=225 30s<=225-100=125 s<=125/30 =4.33, so can buy a maximum of 4 sweaters